OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5

OpenStudy (anonymous):

A. y – 4 = 5(x – 2) B. y – 4 = 5(x + 2) C. y + 4 = 5(x + 2) D. y + 4 = 5x – 10

OpenStudy (anonymous):

Formula is \[(y1-y2)= m (x1-x2)\]

OpenStudy (anonymous):

Plug the Y value into Y2 and the X value into X2 and solve.

OpenStudy (anonymous):

\[(y1-4)=m(x1+2) ?\]